# SORSA: Singular Values and Orthonormal Regularized Singular Vectors Adaptation

### v3

@@ Abstract
The rapid advancement in large language models (LLMs) comes with a significant increase in their parameter size, presenting challenges for adaptation and fine-tuning. Parameter-efficient fine-tuning (PEFT) methods are widely used to adapt LLMs for downstream tasks efficiently. In this paper, we propose Singular Values and Orthonormal Regularized Singular Vectors Adaptation, or SORSA, a novel PEFT method. We introduce a method to analyze the variation of the parameters by performing singular value decomposition (SVD) and discuss and analyze SORSA’s superiority in minimizing the alteration in the SVD aspect. Each SORSA adapter consists of two main parts: trainable principal singular weights $W_{p}=U_{p}\Sigma_{p}V^{\top}_{p}$, and frozen residual weights $W_{r}=U_{r}\Sigma_{r}V^{\top}_{r}$. These parts are initialized by performing SVD on pre-trained weights. Moreover, we implement and analyze an orthonormal regularizer, which we prove could decrease the condition number of $W_{p}$ and allows the optimization to be more efficient. SORSA adapters could be merged during inference, thus eliminating any inference latency. After all, SORSA shows a faster convergence than PiSSA and LoRA in our experiments. On the MATH benchmark, Llama 2 7B adapted using SORSA achieved 10.36% accuracy, outperforming LoRA (5.50%), Full FT (7.22%), and PiSSA (7.44%). On the GSM-8K benchmark, SORSA achieved 56.03% accuracy, surpassing LoRA (42.30%), Full FT (49.05%), and PiSSA (53.07%). We conclude that SORSA offers a new perspective on parameter-efficient fine-tuning, demonstrating remarkable performance. The code is available at https://github.com/Gunale0926/SORSA.
Figure 1: Architecture of a SORSA adapter. We only train parts rendered in orange ($U_{p}$, $\Sigma_{p}$ and $V^{\top}_{p}$), and freeze parts rendered in blue ($U_{r}$, $\Sigma_{r}$ and $V^{\top}_{r}$).
@@ 1 Introduction
Pre-trained large language models (LLMs) show remarkable generalization abilities, allowing them to perform various kinds of natural language processing (NLP) tasks. For specific downstream tasks, full parameter fine-tuning, which continues training all parameters of LLMs on downstream data, is widely used.
However, as the number of parameters in LLMs rapidly increases, full parameter fine-tuning becomes increasingly inefficient. For example, the estimated VRAM requirement for fully fine-tuning Llama 2 7B using Float32 could approach approximately 100 GB, making it unlikely to fully fine-tune the model on a single GPU with current technology. Additionally, the VRAM requirement for fully fine-tuning Llama 2 70B using Float32 exceeds 1 TB, thus rendering it unfeasible on a single GPU with current technology.
To address these challenges, several parameter-efficient fine-tuning (PEFT) methods (Houlsby et al., 2019; Lester et al., 2021; Hu et al., 2021) have been proposed. These methods enable the training of only a limited number of parameters, which significantly reduces VRAM requirements while achieving comparable or even superior performance to full fine-tuning. For instance, tuning Llama 2 7B in Float32 by LoRA with a rank of 128 only takes approximately 60GB VRAM, which allows training in 1 $\times$ NVIDIA A100-80GB or 3 $\times$ NVIDIA RTX 4090-24GB.
@@
In this paper, we propose a novel PEFT approach, Singular Values and Orthonormal Regularized Singular Vectors Adaptation, or SORSA. A SORSA adapter has two main parts: principal singular weights $W_{p}=U_{p}\Sigma_{p}V^{\top}_{p}$, and residual weights $W_{r}=U_{r}\Sigma_{r}V^{\top}_{r}$. These two parts are initialized by performing singular value decomposition (SVD) on pre-trained weight. Residual singular values and vectors will be merged into one matrix and frozen while training. We only train principal singular values and vectors with an orthonormal regularizer implemented to keep the orthonormality of $U_{p}$ and $V_{p}^{\top}$. The architecture of a SORSA adapter is illustrated in Figure 1.
@@
In our experiments, SORSA retains all the benefits of LoRA and its variants while demonstrating remarkable performance compared to PiSSA, LoRA, and full parameter fine-tuning.
@@ 2 Related Works
LoRA (Hu et al., 2021) and its variants are the most popular type of PEFT methods. This type of PEFT is popular for its on-par or better performance than full parameter fine-tuning without introducing any inference latency. LoRA could be represented by equation $W=W_{0}+BA$, where $W_{0}\in\mathbb{R}^{m\times n}$ is the pre-trained weight, $A\in\mathbb{R}^{m\times r}$, using Gaussian initialization, and $B\in\mathbb{R}^{r\times n}$, using zero initialization, are low-rank matrices. Its variant, for example, AdaLoRA (Zhang et al., 2023) introduces an SVD decomposition and pruning for least significant singular values for more efficient parameter updating. DoRA (Liu et al., 2024) proposed a novel way to decompose weight into direction and magnitude by $W=\underline{m}\frac{W_{0}+\underline{BA}}{\|W_{0}+\underline{BA}\|_{c}}$, where $\underline{m}$ is initialized by $\underline{m}=\|W_{0}+\underline{BA}\|_{c}$, $\|\cdot\|_{c}$ denotes column-wise norm. The results show that DoRA has a better learning capacity than LoRA. However, DoRA introduced a calculation of norm in every training step, which makes it much more inefficient in training than LoRA. PiSSA (Meng et al., 2024) decomposes pre-trained weight $W_{0}=U\text{diag}(S)V^{\top}$ by Singular Value Decomposition (SVD), and then splits $W_{0}$ into $W_{pri}$ and $W_{res}$: $W_{pri}=BA$ which is trainable, where $A=U_{[:r]}\text{diag}(S^{\frac{1}{2}}_{[:r]})$ and $B=\text{diag}(S^{\frac{1}{2}}_{[:r]})V^{\top}_{[:r]}$; $W_{res}=U_{[:r,:r]}\text{diag}(S_{[:r]})V^{\top}_{[:r,:]}$ which is frozen. PiSSA results a faster convergence speed and better fitting compared to LoRA. SORSA has a similar architecture to PiSSA, which both conducts SVD and replaces pre-trained weights with residual singular weights. SORSA inherits LoRA and its variants’ benefits, including low training VRAM requirement, no inference burden, and versatility in different architectures.
@@ 3.1 Singular Value Decomposition
According to the meaning of SVD, giving a matrix $W\in\mathbb{R}^{m\times n}$, let $k=\min(m,n)$, we could perform SVD to decompose $W$ by $W=U\Sigma V^{\top}$. Here, $U\in\mathbb{R}^{m\times k}$ are left singular vectors and have orthonormal columns, $V^{\top}\in\mathbb{R}^{k\times n}$ are right singular vectors and have orthonormal rows, and $\Sigma\in\mathbb{R}^{k\times k}$ is a diagonal matrix, where diagonal values are singular values $\sigma^{1},\sigma^{2}\ldots\sigma^{k}$ arranged in descending order.
@@ 3.2 Analysis Method
The singular value and vector variations between pre-trained weight $W_{0}\in\mathbb{R}^{m,n}$ and tuned weight $W_{t}\in\mathbb{R}^{m,n}$, which $t$ denotes the training step, could be defined as follows:
@@
We adopt the analysis on Llama 2 7B (Touvron et al., 2023) using the first 100K data of MetaMathQA (Yu et al., 2024). We test partial fine-tuning, LoRA, and SORSA (with regularizer and without regularizer). See Appendix A.1 for training details of the analysis.
@@ 3.3 Analysis Result
This section analyzes the results of different training methods: partial fine-tuning, LoRA, and SORSA based on the data we collect. The analysis data is illustrated in Figure 2. The four methods obtained similar final loss ($\pm 10\%$).
@@
4 Singular Value and Orthonormal Regularized Singular Vector Adaptation
@@
$W_{p}=U_{[:,:r]}\Sigma_{[:r,:r]}V^{\top}_{[:r,:]}\in\mathbb{R}^{m\times n}$ (5)
$W_{r}=U_{[:,r:]}\Sigma_{[r:,r:]}V^{\top}_{[r:,:]}\in\mathbb{R}^{m\times n}$ (6)
Here, $U$ represents the left singular vectors, $\Sigma$ represents the diagonal matrix of singular values, and $V^{\top}$ represents the right singular vectors. We use PyTorch (Paszke et al., 2019) syntax to demonstrate matrix selection, where $[:,:r]$ denotes selecting the first $r$ columns of the matrix, and $[r:,:]$ denotes selecting the last $r$ rows of the matrix. We rewrite $U_{[:,:r]}$, $\Sigma_{[:r,:r]}$ and $V^{\top}_{[:r,:]}$, which consist $W_{p}$ as $U_{p}$, $\Sigma_{p}$ and $V^{\top}_{p}$ for simplicity, and rewrite $U_{[:,r:]}$, $\Sigma_{[r:,r:]}$ and $V^{\top}_{[r:,:]}$, which consist $W_{r}$ as $U_{r}$, $\Sigma_{r}$ and $V^{\top}_{r}$ correspondingly.
The initialization of $W_{r}$ in SORSA is the same as PiSSA (Meng et al., 2024). Nevertheless, unlike PiSSA which merge $\Sigma_{p}$ with $U_{p}$ and $V^{\top}_{p}$ into $A$ and $B$ by $A=U_{p}\Sigma_{p}^{\frac{1}{2}}$ and $B=\Sigma_{p}^{\frac{1}{2}}V^{\top}_{p}$, SORSA remains $U_{p}$, $\Sigma_{p}$, and $V^{\top}_{p}$ in separate matrices. SORSA is defined by Eq equation 7, which is initially equivalent to the pre-trained weight $W_{0}$.
During training, $W_{r}$ remains frozen, and only $U_{p}$, $\Sigma_{p}$, and $V^{\top}_{p}$ are updated. Since $\Sigma_{p}$ is a diagonal matrix, only the diagonal elements are updated, while the rest of the matrix remains zero.
$\text{SORSA}(x):=W_{r}x+W_{p}x$ (7)
@@
Therefore, parameter updating of $U_{p}$, $\Sigma_{p}$ and $V^{\top}_{p}$ in a SORSA adapter at training step $t$ could be expressed as:
@@
We update the SORSA as Eq equation 10 for implementation simplicity.
@@ 5 Gradient Analysis
The proofs of the theorems and lemmas, along with additional mathematical details, are provided in Appendix B.
@@
$\gamma<\frac{2}{k\eta_{d}}-L_{train}$ (11)
@@
Lemma 5.4.
Let $W_{p}^{\text{unreg}}=U_{p}^{\text{unreg}}\Sigma_{p}^{\text{unreg}}(V_{p}^{\text{unreg}})^{\top}$ be the $W_{p}$ only training without using regularizer, and $W_{p}^{\text{reg}}=U_{p}^{\text{reg}}\Sigma_{p}^{\text{reg}}(V_{p}^{\text{reg}})^{\top}$ be the $W_{p}$ training with the regularizer. For each singular value $\sigma_{i}$, the following bound holds:
@@
Lemma 5.4 provides a crucial connection between the regularizer and the singular values. Building on this result, we arrive at our main theorem regarding the condition number:
Theorem 5.5.
@@
When training ($t>0$), while $U_{p}^{\text{unreg}}$ and $(V_{p}^{\text{unreg}})^{\top}$ are more and more deviating from orthonormal, $\kappa(U_{p,\text{unreg}})\cdot\kappa(V_{p,\text{unreg}}^{\top})$ will increase, and $(1+\delta_{1,t})(1+\delta_{2,t})(\frac{1+\epsilon_{t}}{1-\epsilon_{t}})$ will generally remain $1$ due to the regularization in orthonormality of $U_{p}^{\text{reg}}$ and $(V_{p}^{\text{reg}})^{\top}$. When $\kappa(U_{p,\text{unreg}})\cdot\kappa(V_{p,\text{unreg}}^{\top})$ becomes greater than $(1+\delta_{1,t})(1+\delta_{2,t})(\frac{1+\epsilon_{t}}{1-\epsilon_{t}})$, the ratio becomes strictly less than 1, indicating an improvement in the condition number.

### v4

@@ Abstract
The rapid advancement in large language models (LLMs) comes with a significant increase in their parameter size, presenting challenges for adaptation and fine-tuning. Parameter-efficient fine-tuning (PEFT) methods are widely used to adapt LLMs for downstream tasks efficiently. In this paper, we propose Singular Values and Orthonormal Regularized Singular Vectors Adaptation, or SORSA, a novel PEFT method. We introduce a method to analyze the variation of the parameters by performing singular value decomposition (SVD) and discuss and analyze SORSA’s superiority in minimizing the alteration in the SVD aspect. Each SORSA adapter consists of two main parts: trainable principal singular weights $W_{p}=U_{p}\text{diag}(S_{p})V^{\top}_{p}$, and frozen residual weights $W_{r}=U_{r}\text{diag}(S_{r})V^{\top}_{r}$. These parts are initialized by performing SVD on pre-trained weights. Moreover, we implement and analyze an orthonormal regularizer, which we prove could decrease the condition number of $W_{p}$ and allows the optimization to be more efficient. SORSA adapters could be merged during inference, thus eliminating any inference latency. After all, SORSA shows a faster convergence than PiSSA and LoRA in our experiments. On the GSM-8K benchmark, Llama 2 7B adapted using SORSA achieved 56.03% accuracy, surpassing LoRA (42.30%), Full FT (49.05%), and PiSSA (53.07%). On the MATH benchmark, SORSA achieved 10.36% accuracy, outperforming LoRA (5.50%), Full FT (7.22%), and PiSSA (7.44%). We conclude that SORSA offers a new perspective on parameter-efficient fine-tuning, demonstrating remarkable performance. The code is available at https://github.com/Gunale0926/SORSA.
Figure 1: Architecture of a SORSA adapter. We only train parts rendered in orange ($U_{p}$, $\text{diag}(S_{p})$ and $V^{\top}_{p}$), and freeze parts rendered in blue ($U_{r}$, $\text{diag}(S_{r})$ and $V^{\top}_{r}$).
@@ 1 Introduction
Pre-trained large language models (LLMs) show remarkable generalization abilities, allowing them to perform various kinds of natural language processing (NLP) tasks (Peng et al., 2024; Touvron et al., 2023; Dubey et al., 2024; Radford et al., 2019; OpenAI, 2023). For specific downstream tasks, full parameter fine-tuning, which continues training all parameters of LLMs on downstream data, is widely used.
However, as the number of parameters in LLMs rapidly increases, full parameter fine-tuning becomes increasingly inefficient. For example, the estimated VRAM requirement for fully fine-tuning Llama 2 7B using Float32 could approach approximately 100 GB, making it unlikely to fully fine-tune the model on a single GPU with current technology. Additionally, the VRAM requirement for fully fine-tuning Llama 2 70B using Float32 exceeds 1 TB (Touvron et al., 2023; Anthony et al., 2023), thus rendering it unfeasible on a single GPU with current technology.
To address these challenges, several parameter-efficient fine-tuning (PEFT) methods (Houlsby et al., 2019; Lester et al., 2021; Hu et al., 2021) have been proposed. These methods enable the training of only a limited number of parameters, which significantly reduces VRAM requirements while achieving comparable or even superior performance to full fine-tuning. For instance, tuning Llama 2 7B in Float32 by LoRA (Hu et al., 2021) with a rank of 128 only takes approximately 60GB VRAM, which allows training on 1 $\times$ NVIDIA A100 (80GB), or even 3 $\times$ NVIDIA RTX 4090 (24GB).
@@
In this paper, we propose a novel PEFT approach, Singular Values and Orthonormal Regularized Singular Vectors Adaptation, or SORSA. A SORSA adapter has two main parts: principal singular weights $W_{p}=U_{p}\text{diag}(S_{p})V^{\top}_{p}$, and residual weights $W_{r}=U_{r}\text{diag}(S_{r})V^{\top}_{r}$. These two parts are initialized by performing singular value decomposition (SVD) on pre-trained weight. Residual singular values and vectors will be merged into one matrix and frozen while training. We only train principal singular values and vectors with an orthonormal regularizer implemented to keep the orthonormality of $U_{p}$ and $V_{p}^{\top}$. The architecture of a SORSA adapter is illustrated in Figure 1.
@@
SORSA retains all the benefits of LoRA and its variants while demonstrating remarkable performance compared to PiSSA, LoRA, and full parameter fine-tuning in our experiments.
@@ 2 Related Works
LoRA (Hu et al., 2021) and its variants are the most popular type of PEFT methods. This type of PEFT is popular for its on-par or better performance than full parameter fine-tuning without introducing any inference latency. LoRA could be represented by equation $W=W_{0}+BA$, where $W_{0}\in\mathbb{R}^{m\times n}$ is the pre-trained weight, $A\in\mathbb{R}^{m\times r}$, using Gaussian initialization, and $B\in\mathbb{R}^{r\times n}$, using zero initialization, are low-rank matrices. Its variant, for example, AdaLoRA (Zhang et al., 2023) introduces an SVD decomposition and pruning for least significant singular values for more efficient parameter updating. DoRA (Liu et al., 2024) proposed a novel way to decompose weight into direction and magnitude by $W=\underline{m}\frac{W_{0}+\underline{BA}}{\|W_{0}+\underline{BA}\|_{c}}$, where $\underline{m}$ is initialized by $\underline{m}=\|W_{0}+\underline{BA}\|_{c}$, $\|\cdot\|_{c}$ denotes column-wise norm. The results show that DoRA has a better learning capacity than LoRA. However, DoRA introduced a calculation of norm in every training step, which makes it much more inefficient in training than LoRA. OLoRA (Büyükakyüz, 2024) uses QR decomposition to initialize the LoRA adapters $A$ and $B$, which initializes $B$ as an orthogonal matrix. They discusses the significance of orthonormality in neural networks’ weight (See Section 5 for more details). In their experiments, OLoRA demonstrates faster convergence than LoRA. PiSSA (Meng et al., 2024) decomposes pre-trained weight $W_{0}=U\text{diag}(S)V^{\top}$ by Singular Value Decomposition (SVD), and then splits $W_{0}$ into $W_{pri}$ and $W_{res}$: $W_{pri}=AB$ which is trainable, where $A=U_{[:r]}\text{diag}(S^{\frac{1}{2}}_{[:r]})$ and $B=\text{diag}(S^{\frac{1}{2}}_{[:r]})V^{\top}_{[:r]}$; $W_{res}=U_{[:r,:r]}\text{diag}(S_{[:r]})V^{\top}_{[:r,:]}$ which is frozen. PiSSA results a faster convergence speed and better fitting compared to LoRA. SORSA has a similar architecture to PiSSA, which both conducts SVD and replaces pre-trained weights with residual singular weights. SORSA inherits LoRA and its variants’ benefits, including low training VRAM requirement, no inference burden, and versatility in different architectures.
@@ 3.1 Singular Value Decomposition
According to the meaning of SVD, giving a matrix $W\in\mathbb{R}^{m\times n}$, let $k=\min(m,n)$, we could perform SVD to decompose $W$ by $W=U\text{diag}(S)V^{\top}$. Here, $U\in\mathbb{R}^{m\times k}$ is matrix of left singular vectors and have orthonormal columns, $V\in\mathbb{R}^{n\times k}$ is matrix of right singular vectors and have orthonormal columns, and $S\in\mathbb{R}^{k}$ are singular values $\sigma^{1},\sigma^{2}\ldots\sigma^{k}$ arranged in descending order. $\text{diag}(W)$ denotes a function to form a diagonal matrix from $W$.
@@ 3.2 Analysis Method
The singular value and vector variations between pre-trained weight $W_{0}\in\mathbb{R}^{m\times n}$ and tuned weight $W_{t}\in\mathbb{R}^{m\times n}$, which $t$ denotes the training step, could be defined as follows:
@@
We adopt the analysis on Llama 2 7B (Touvron et al., 2023) using the first 100K data of MetaMathQA (Yu et al., 2024). We test partial fine-tuning, LoRA, and SORSA (with regularizer and without regularizer). See Section B.1 for training details of the analysis.
@@ 3.3 Analysis Result
This section analyzes the results of different training methods: partial fine-tuning, LoRA, and SORSA based on the data we collect. The analysis data is illustrated in Figure 2.
@@
4 SORSA: Singular Value and Orthonormal Regularized Singular Vector Adaptation
@@
$W_{p}=U_{[:,:r]}\text{diag}(S_{[:r]})V^{\top}_{[:r,:]}\in\mathbb{R}^{m\times n}$ (5)
$W_{r}=U_{[:,r:]}\text{diag}(S_{[r:]})V^{\top}_{[r:,:]}\in\mathbb{R}^{m\times n}$ (6)
Here, $U$ represents the matrix of left singular vectors, $S$ represents the singular values, $\text{diag}(W)$ denotes a function to form a diagonal matrix from $W$, and $V$ represents the matrix of right singular vectors. We use PyTorch (Paszke et al., 2019) syntax to demonstrate matrix selection, where $[:,:r]$ denotes selecting the first $r$ columns of the matrix, and $[r:,:]$ denotes selecting the last $r$ rows of the matrix. We rewrite $U_{[:,:r]}$, $S_{[:r]}$ and $V^{\top}_{[:r,:]}$, which consist $W_{p}$ as $U_{p}$, $S_{p}$ and $V^{\top}_{p}$ for simplicity, and rewrite $U_{[:,r:]}$, $S_{[:r]}$ and $V^{\top}_{[r:,:]}$, which consist $W_{r}$ as $U_{r}$, $S_{r}$ and $V^{\top}_{r}$ correspondingly.
The initialization of $W_{r}$ in SORSA is the same as PiSSA (Meng et al., 2024). Nevertheless, unlike PiSSA which merge $\text{diag}(S_{p})$ with $U_{p}$ and $V^{\top}_{p}$ into $A$ and $B$ by $A=U_{p}\text{diag}(S_{p})^{\frac{1}{2}}$ and $B=\text{diag}(S_{p})^{\frac{1}{2}}V^{\top}_{p}$, SORSA remains $U_{p}$, $S_{p}$, and $V^{\top}_{p}$ in separate matrices. SORSA is defined by Equation 7, which is initially equivalent to the pre-trained weight $W_{0}$.
During training, $W_{r}$ remains frozen, and only $U_{p}$, $S_{p}$, and $V^{\top}_{p}$ are updated.
SORSA is defined as:
$\text{SORSA}(x):=x(W_{r}+W_{p})=xW_{r}+xU_{p}\text{diag}(S_{p})V^{\top}_{p}$ (7)
In our implementation, we use a optimized version of SORSA equation which results a much faster computation speed, and could approach to the speed of LoRA. See Appendix A for more details.
@@
Therefore, parameter updating of $W_{p}$ in a SORSA adapter at training step $t$ could be expressed as:
@@
We update the SORSA as Equation 10 for implementation simplicity.
@@ 5 Gradient Analysis
The proofs of the theorems and lemmas, along with additional mathematical details, are provided in Appendix C.
@@
$0\leq\gamma<\frac{2}{k\eta_{d}}-L_{train}$ (11)
@@
Lemma 4.
Let $W_{p}^{\text{unreg}}=U_{p}^{\text{unreg}}\text{diag}(S_{p})^{\text{unreg}}(V_{p}^{\text{unreg}})^{\top}$ be the $W_{p}$ only training without using regularizer, and $W_{p}^{\text{reg}}=U_{p}^{\text{reg}}\text{diag}(S_{p})^{\text{reg}}(V_{p}^{\text{reg}})^{\top}$ be the $W_{p}$ training with the regularizer. For each singular value $\sigma_{i}$, the following bound holds:
@@
Lemma 4 provides a crucial connection between the regularizer and the singular values. Building on this result, we arrive at our main theorem regarding the condition number:
Theorem 5.
@@
When training ($t>0$), while $U_{p}^{\text{unreg}}$ and $(V_{p}^{\text{unreg}})^{\top}$ are more and more deviating from orthonormal, $\kappa(U_{p}^{\text{unreg}})\cdot\kappa((V_{p}^{\text{unreg}})^{\top})$ will increase, and $(1+\delta_{1,t})(1+\delta_{2,t})(\frac{1+\epsilon_{t}}{1-\epsilon_{t}})$ will generally remain $1$ due to the regularization in orthonormality of $U_{p}^{\text{reg}}$ and $(V_{p}^{\text{reg}})^{\top}$. When $\kappa(U_{p}^{\text{unreg}})\cdot\kappa((V_{p}^{\text{unreg}})^{\top})$ becomes greater than $(1+\delta_{1,t})(1+\delta_{2,t})(\frac{1+\epsilon_{t}}{1-\epsilon_{t}})$, the ratio becomes strictly less than 1, indicating an improvement in the condition number.
@@
Moreover, as mentioned in Büyükakyüz (2024), orthonormal matrices in neuron networks could improve gradient flow (Saxe et al., 2014; Arjovsky et al., 2016) and enhanced optimization landscape (Huang et al., 2018; Wisdom et al., 2016), which could also explain SORSA’s superior performance in convergence.